To be honest I don't know what kind of problem this is. Its extra credit if we can answer it but I have no clue, can u help please?

3x + 7 = x^2 + k = 7x + 15 What is the value of k?

Answer by fractalier(6550)

Well if
3x + 7 = 7x + 15
we can find x...
-8 = 4x and
x = -2
now plug in
3(-2) + 7 = (-2) + k
-6 + 7 = k - 2
1 = k - 2
k = 3